. An alloy of copper , gold and silver is found to have Cu atoms forming simple cubic lattice . If Ag atoms occupy all the face centres and Au atoms at body centres what is the simplest formula of the alloy?

Grade:12

1 Answers

Vikas TU
7 years ago
Cu = 8*1/8 = 1 (SCC)
Ag = 6*1/2 = 3 (FCC)
Au = 1(Body Centre)
Hence the formulae would become:
CuAg3Au